# “Did They Consent to That?”: Safer Digital Intimacy via Proactive Protection Against Image-Based Sexual Abuse

### v1

@@ Abstract
Trigger warning: This work addresses sexual violence.
@@ 1 Introduction
Sharing intimate content [^1] We define intimate content as images that: show a nude or semi-nude subject, contain intimate body parts, and/or intend to arouse. Throughout the text of this paper, we use the term images to refer to photos and videos. is a common practice. While prevalence is difficult to measure, estimates suggest that as many as 8 in every 10 adults share intimate content [78, 24, 73]. Sharing intimate content can benefit relationship intimacy [48, 29, 81, 6, 52, 73], body positivity [81, 24, 34], and offer employment [34, 67, 69].
Societal norms and a lack of technological mitigations put people at risk of experiencing sexual abuse when they share intimate content. Image-based sexual abuse (IBSA) includes a spectrum of abuse from the nonconsensual distribution of consensually-created intimate content (NDII) [64] and the threat of NDII, to the nonconsensual creation of intimate imagery (e.g., via hidden cameras, “upskirting,” extortion, or AI) [55, 40, 23, 50] or the unsolicited sharing of intimate content[5, 22]. Studies estimate that one-third of individuals have been subjected to IBSA [60] (similar to rates of other types of sexual violence [12]). However, it is difficult to accurately estimate prevalence, as victim-survivors [^2] We use the terminology victim-survivor to capture the range of ways in which people who have experienced IBSA may identify [87, 72, 40]. face significant stigma in reporting experiences and may not always be aware that content has been created or shared without their consent [40].
@@
We conclude in Section 9 by synthesizing the results of our study with the state-of-the-art to offer a set of concrete recommendations to platforms and S&P researchers on next steps to prevent NDII.
@@ 2 Related Work
Technology & Intimate Content. Work exploring this topic from a technological viewpoint is scarce [41]. Notably, Geeng et al. surveyed 247 U.S. adults quantifying their general sexting practices and security concerns [29]. Respondents’ sexting concerns included regret after sharing content, unwanted sexual content, NDII, and data breaches. The authors found that their respondents managed their concerns by selecting platforms based on availability of specific features (e.g. expiring messages). Additionally, they found evidence of behavioral strategies used to fill in the gaps left by platforms: respondents reported limiting what they share with whom and establishing rules with recipients.
Relatedly, Waldman studied the behavioral mitigation strategies used by “917 mostly gay male and bisexual users of geosocial dating apps” and found that participants faced an implicit expectation to share intimate content in participating on the apps  [81]. To mitigate risks, participants would delay sending images, avoid identifying features from photos, or select photos that they would be less embarrassed by in case of NDII. Finally, examining trust in platform features that accomodate sexting use cases, Roesner et al. found that despite media focus on Snapchat’s expiring messages feature for sexting, less than 2% of their survey respondents used Snapchat primarily for sexting and fewer than 15% used the app for sexting at all [61] because they did not trust Snapchat nor its security practices and features.
We expand on these works, which provide a quantitative analysis of people’s sexting practices and concerns, by using qualitative methods to focus on how people use specific strategies to defend against NDII-related threats and why and what gaps remain between their threat models and the defensive strategies available to them. In so doing, we answer Geeng et al.’s call for future work to “identify—and bridge—any gaps between the threat models [platforms] securely support and the threat models important to individuals engaging in sexting” by systematically studying the “technological properties” offered by platforms; we additionally answer the call of criminology scholarship on NDII to specifically use qualitative methods to understand use of and potential for technological mitigation in this complex and nuanced space [39].
Finally, multiple works have explored user perceptions of security in messaging apps (often used for sharing intimate content) [2, 3, 21, 58, 37, 1, 82]. They found that users had large misconceptions from inconsistencies in (or lack of) information on end-to-end encryption (E2EE). In lieu of reliable information, users may turn to folk theories to inform their decision-making [85] and develop a false sense of security or blanket mistrust. Our findings are consistent with these results and we expand on this literature by including user perceptions of secure messaging in this context.
@@
IBSA is a wrong in and of itself: “regardless of the nature and extent of any further consequential harm…[IBSA] constitutes [a] breach of an individual’s 'fundamental rights to dignity and privacy, as well as their freedom of sexual expression and autonomy'” [54]. In the largest qualitative study of IBSA victim-surivivors to date ($n=75$), McGlynn et al. describe IBSA as “torture for the soul," specifically noting the devastating disruption to participants’ lives, work, and selves as well as the permanence and “endlessness” of the abuse and its impacts [54].
Beyond the wrong of IBSA alone, victim-survivors can experience harmful mental health consequences ranging from high levels of stress to suicidal thoughts and diagnoses of post-traumatic stress disorder, depression, and anxiety [8, 42, 22]. IBSA victim-survivors are often subjected to further harms as they navigate the post-IBSA landscape by seeking justice, advice, legal assistance, and mental health support. A survey among 6,109 individuals from New Zealand, Australia, and the UK found that one-third of respondents expressed “victim-blaming sentiments," defined as attributing blame towards victims and suggesting that they are at least partially responsible if an intimate image ends up online [25]. Victim-blaming has been associated with poorer outcomes for victim-survivors, acting as a barrier for reporting and help-seeking, worsening victim-survivors’ already impacted mental health [57] [59]. Accordingly, approximately 73% of IBSA victim-survivors choose not to report out of fear or embarrassment [63].
@@
Legal Perspectives. More laws are being introduced globally to address NDII [18, 79, 30, 75, 45]. While legislation is an important avenue for justice and changing societal perceptions [14], legal scholars have described its many limitations. Legislation may evolve too slowly to keep up with ever changing digital threats or be too narrow to address international platforms [14]. Litigation is expensive and victim-survivors may not know who perpetrated the abuse, may fear retaliation or being in a courtroom with their abuser, or may not want to reveal their personal information to courts [13]. As a result, it is critical to explore interventions on the technological side, especially those that platforms can implement [39, 56].
@@ 3 Methodology
We conducted 52 semi-structured interviews (August-November 2022) with adults (over 18 years old) living in European countries [89] who had shared intimate content in the past year: 24 with sex workers who produced their own digital content and 28 with those who shared for noncommercial purposes. 42% of all participants were victim-survivors of IBSA (predominantly NDII) [^5] One participant had experienced the nonconsensual creation of intimate content..
Recruitment. We recruited participants through convenience sampling via social media, relevant community organizations and personal contacts. We detail this process in Appendix C.1. To capture diverse perspectives, all participants were asked questions in a screening survey about their education, industries of employment, and age. They could optionally respond if they have a disability and self-describe their gender, race, and sexual orientation. Participants were compensated €50 (or equivalent amount) through PayPal or Amazon giftcard.
@@
Interview Data Collection. Four of the authors conducted the semi-structured interviews. Sex-working participants were asked about their commercial and recreational content, if applicable. The interview protocols included several topics related to sharing intimate content (see Appendix C.4). We asked about the context in which participants shared intimate content (e.g. with whom, their motivations) and the technologies they used for storing and sharing intimate content. Then we asked questions on how they navigate trust with recipients, security and safety concerns, and proactive practices. Based on existing literature and media advice on sharing intimate content discussed in Section 2, we selected a set of features that we explicitly asked about (listed in Appendix C.4). Lastly, we asked participants to imagine potential futures[35] for design changes and proposals for better technologies for intimate content. Although we asked victim-survivors about their experiences with NDII, the focus of this work is on proactive mitigation. We will analyze victim-survivor narratives on post-harm mitigation in follow-up work.
@@
Analysis. We followed an open coding analysis process. We randomly selected 4 interviews (2 commercial and 2 recreational) that three authors independently coded. They each created a codebook of main themes based on these interviews. Four of the authors (including the coders) then met to create the codebook. As part of this process, we grouped smaller codes into larger themes. The coders then each randomly selected 4 interviews to code using the developed codebook. They reached “substantial” intercoder agreement as measured by Krippendorf’s Alpha [47] and met to resolve any disagreements. They collectively finalized the codebook and then divided and independently coded the transcripts. Throughout the paper we provide a relative sense of the frequency of the themes we coded (e.g., that a theme was mentioned by around a quarter of participants) but emphasize the importance of not reducing “sensitivity to complex concepts and nuances in data” by viewing these data through a quantitative lens [53].
Demographics. Participants reported that they resided in 15 different European countries. 65% of our participants identified as White, 35% did not. The majority (71%) self-described as queer, asexual, bisexual, pansexual, gay, or lesbian. 73% held at least a Bachelor’s degree and 56% responded that they had a disability. 27% were cis or trans women, 38% were cis or trans men, 19% were non-binary, gender fluid, or another gender. Appendix C.3 contains additional details.
@@
Ethics. This work was approved by our institution’s ethics review board. We took great care to protect the privacy of participants, which we detail in Appendix C.2. Since 42% of the participants had experienced NDII, we warned all participants that discussing their experiences may be triggering and phrased questions to avoid re-traumatization. To ensure both participant and researcher protection, our team includes a clinical psychologist who reviewed the interview protocols, analyzed data, and supported the research team. We additionally asked our partner organization, the European Sex Workers Rights Alliance (ESWA), and a victim-survivor of NDII for feedback on the protocol. We followed trauma-informed best practices set forth and followed by sexual violence researchers [11] that minimize risk to participants and center victim-survivors’ choice, well-being, empowerment, and autonomy. Team members were trained in response techniques by a licensed clinical psychologist who specializes in sexual assault and trauma (e.g. offering to pause/stop the interview) and were prepared with resources (e.g., hotline contact information) to offer participants if needed. We note that sharing these experiences may also be a form of empowerment for victim-survivors [32].
@@
Impacts and Research Justice. The interviews were transcribed by a sex worker (trained in transcription and compensated) as part of a research justice commitment to materially support communities we study. We published a white paper of this work through ESWA. We will share published research with participants who requested it and plan to develop a resource guide based on this work. In addition, we developed wireframes that we used to communicate design recommendations to platform product teams and to policymakers. These findings are also being used in IBSA workshops/events hosted by ESWA.
@@
Positionality. This work was conducted by a team of researchers who collectively are scholars of technology, sex work and sexual violence. This research focuses on adults living in Europe but only half of the research team is based in Europe. Given the diversity of our participants, our team does not completely reflect all of the genders, ethnicities, and sexual orientations of our participants. Therefore, there may have been some limitations in our interpretations of the data.
@@
Limitations. Since we conducted interviews in English and it is not the primary language in many European countries, this may have limited or skewed participation to those in countries where English is more prevalent and/or to those who are more educated, who tend to have studied English. Although we re-assured participants that we were interested in understanding their perspectives to advocate for safer technologies for intimate content and took care to phrase questions non-judgmentally, the existing stigma of sharing intimate content may have prevented participants from fully disclosing their behaviors. Since we asked about specific safety features, there may be more data for them but for each of these features, there were participants who independently mentioned them.
@@
Platform Blinding. With the exception of a few quotes, we replace specific platforms with the type of platform mentioned in italics. A quote that says, “I use WhatsApp” would be replaced as, “I use [messaging app].” This is to draw focus away from specific platforms and emphasize that intimate content is likely being shared across any and all platforms that enable the storing and/or sharing of images or videos.
@@ 4 Summary of Findings
We conducted an interview study with 52 people in Europe who share intimate content (24 sex workers, 28 recreational sharers), 42% of whom were victim-survivors of IBSA. Our participants used 40+ technologies to share intimate content. They relied on a combination of heuristics to choose technologies for sharing but nearly half chose to use their default method of communication. Participants’ threat models around NDII were focused on a recipient resharing their content but they were also concerned about data breaches, hacking, and unintentional content discovery. Echoing prior research (Section 2), all emphasized that stigma exacerbated these threats. To protect themselves, participants implemented a maze of safety strategies that included technical features (e.g. screenshot notifications) and interpersonal strategies (e.g. setting rules). In sections 5 through 8, we detail sharing contexts, threat models, how participants chose technologies for sharing content, and defensive strategies. In Section 9, we include narratives on how technology could be improved to support digital intimacy/reduce NDII alongside recommendations for platforms and researchers.

### v2

# “Did They F***ing Consent to That?”: Safer Digital Intimacy via Proactive Protection Against Image-Based Sexual Abuse

@@ Abstract
Content Warning: This work discusses sexual violence, specifically, the harms of image-based sexual abuse (particularly in Sections  2 and  6).
@@ 1 Introduction
Sharing intimate content [^1] We define intimate content as images that: show a nude or semi-nude subject, contain intimate body parts, and/or intend to arouse. Throughout the text of this paper, we use the term images to refer to photos and videos. is a common practice. While prevalence is difficult to measure, estimates suggest that as many as 8 in every 10 adults share intimate content [78, 24, 73]. Sharing intimate content can benefit relationship intimacy [48, 29, 81, 6, 52, 73], body positivity [81, 24, 34, bianchi_individual_2021], and offer employment [34, 67, 69].
Societal norms and a lack of technological mitigations put people at risk of experiencing sexual violence when they share intimate content. Image-based sexual abuse (IBSA) encompasses a spectrum of violence, including (1) the nonconsensual distribution of consensually-created intimate content (NDII) [64] and the threat of NDII, (2) the nonconsensual creation of intimate imagery (e.g., via hidden cameras, “upskirting,” extortion, or AI) [55, 40, 23, 50], and (3) the unsolicited sharing of intimate content[5, 22]. Studies estimate that one-third of individuals have been subjected to IBSA [60] (similar to rates of other types of sexual violence [12]). However, it is difficult to accurately estimate prevalence, as victim-survivors [^2] We use the terminology victim-survivor to capture the range of ways in which people who have experienced IBSA may identify [87, 72, 40]. face significant stigma in reporting experiences and may not always be aware that content has been created or shared without their consent [40].
@@
We conclude in Section 9 by synthesizing the results of our study with the state of the art to offer a set of concrete recommendations to platforms and S&P researchers on next steps to prevent NDII.
@@ 2 Related Work
Technology & Intimate Content. Work exploring this topic from a technological viewpoint is scarce [41]. Notably, Geeng et al. surveyed 247 U.S. adults quantifying their general sexting practices and security concerns [29]. Respondents’ sexting concerns included regret after sharing content, unwanted sexual content, NDII, and data breaches. The authors found that their respondents managed their concerns by selecting platforms based on availability of specific features (e.g., expiring messages). Additionally, they found evidence of behavioral strategies used to fill in the gaps left by platforms: respondents reported limiting what they share depending on the recipient and establishing rules with recipients.
Relatedly, Waldman studied the behavioral mitigation strategies used by “917 mostly gay male and bisexual users of geosocial dating apps” and found that participants faced an implicit expectation to share intimate content in participating on the apps  [81]. To mitigate risks, participants would delay sending images, avoid identifying features from photos, or select photos that they would be less embarrassed by in case of NDII. Finally, examining trust in platform features that accommodate sexting use cases, Roesner et al. found that despite media focus on Snapchat’s expiring messages feature for sexting, less than 2% of their survey respondents used Snapchat primarily for sexting and fewer than 15% used the app for sexting at all [61] because they did not trust Snapchat nor its security practices and features.
We expand on these works, which provide a quantitative analysis of people’s sexting practices and concerns, by using qualitative methods to focus on how people use specific strategies to defend against NDII-related threats and why and what gaps remain between their threat models and the defensive strategies available to them. In so doing, we answer Geeng et al.’s call for future work to “identify—and bridge—any gaps between the threat models [that platforms] securely support and the threat models important to individuals engaging in sexting” by systematically studying the “technological properties” offered by platforms; we additionally answer the call of criminology scholarship on NDII to specifically use qualitative methods to understand use of and potential for technological mitigation in this complex and nuanced space [39].
Lastly, multiple works have explored user perceptions of security in messaging apps (often used for sharing intimate content) [2, 3, 21, 58, 37, 1, 82]. They found that users had large misconceptions from inconsistencies in (or lack of) information on end-to-end encryption (E2EE). In lieu of reliable information, users may turn to folk theories to inform their decision-making [85] and develop a false sense of security or blanket mistrust. Our findings are consistent with these results and we expand on this literature by including user perceptions of secure messaging in this context.
@@
IBSA is a wrong in and of itself: “regardless of the nature and extent of any further consequential harm…[IBSA] constitutes [a] breach of an individual’s 'fundamental rights to dignity and privacy, as well as their freedom of sexual expression and autonomy'” [54]. In the largest qualitative study of IBSA victim-surivivors to date ($n=75$), McGlynn et al. describe IBSA as “torture for the soul,” specifically noting the devastating disruption to participants’ lives, work, and selves as well as the permanence and “endlessness” of the abuse and its impacts [54].
Beyond the wrong of IBSA alone, victim-survivors can experience harmful mental health consequences ranging from high levels of stress to suicidal thoughts and diagnoses of post-traumatic stress disorder, depression, and anxiety [8, 42, 22]. IBSA victim-survivors are often subjected to further harms as they navigate the post-IBSA landscape by seeking justice, advice, legal assistance, and mental health support. A survey among 6,109 individuals from New Zealand, Australia, and the UK found that one-third of respondents expressed “victim-blaming sentiments,” defined as attributing blame towards victims and suggesting that they are at least partially responsible if an intimate image ends up online [25]. Victim-blaming has been associated with poorer outcomes for victim-survivors, acting as a barrier for reporting and help-seeking, worsening victim-survivors’ already-impacted mental health [57] [59]. Accordingly, approximately 73% of IBSA victim-survivors choose not to report out of fear or embarrassment [63].
@@
Legal Perspectives. More laws are being introduced globally to address NDII [18, 79, 30, 75, 45]. While legislation is an important avenue for justice and changing societal perceptions [14], legal scholars have described its many limitations. Legislation may evolve too slowly to keep up with ever changing digital threats or be too narrow to address international platforms [14]. Litigation is expensive and victim-survivors may not know who perpetrated the abuse, may fear retaliation, may not want to be in a courtroom with their abuser, or may not want to reveal their personal information to courts [13]. As a result, it is critical to explore interventions on the technological side, especially those that platforms can implement [39, 56].
@@ 3 Methodology
We conducted 52 semi-structured interviews (August-November 2022) with adults (over 18 years old) living in European countries [89] who had consensually shared intimate content in the past year: 24 with sex workers who produced their own digital content and 28 with those who shared for noncommercial purposes. 42% of all participants were victim-survivors of IBSA (predominantly NDII). [^5] One participant had experienced the nonconsensual creation of intimate content.
Recruitment. We recruited participants through convenience sampling via social media, relevant community organizations and personal contacts. To recruit recreational sexters, two authors posted a recruitment graphic to their social media accounts. To recruit sex workers, we used a similar but sex-work specific graphic (see Appendix B.1). We contacted sex worker rights organizations to share it and the graphic was also posted in sex-worker only social media groups. The graphics included sign-up links that led to a Qualtrics survey that screened out individuals who are younger than 18 years old, had not sent intimate content in the past year, and/or were not living in a European country. At the end of the screening survey, eligible participants were asked to review and sign a consent form for the interview study. We interviewed participants until we reached saturation for each group: recreational non-victim-survivors, recreational victim-survivors, commercial non-victim-survivors, commercial victim-survivors. To capture different perspectives, participants were asked questions related to various dimensions of diversity in the screening survey (see Section 3.1). Participants were compensated €50 (or equivalent amount) through PayPal or Amazon giftcard.
@@
Interview Data Collection. Four of the authors conducted the semi-structured interviews. Sex-working participants were asked about their commercial and recreational content, if applicable. The interview protocols included several topics related to sharing intimate content (see Appendix B.3). We asked about the contexts in which participants shared intimate content (e.g., with whom, their motivations) and the technologies they used for storing and sharing intimate content. We then asked questions on how they navigate trust with recipients, security and safety concerns, and their proactive practices. Based on existing literature and media advice on sharing intimate content discussed in Section 2, we selected a set of features that we explicitly asked about (listed in Appendix B.3). Lastly, we asked participants to imagine potential futures[35] for design changes and proposals for better technologies for intimate content. Although we asked victim-survivors about their experiences with NDII, the focus of this work is on proactive mitigation. We will analyze post-harm mitigation in follow-up work.
@@
Analysis. We followed an open coding analysis process. We randomly selected 4 interviews (2 commercial, 2 recreational) that three authors independently coded. They each created a codebook of main themes based on these interviews. Four of the authors (including the coders) then met to create the codebook. As part of this process, we grouped smaller codes into larger themes. The coders then each randomly selected 4 interviews to code using the developed codebook. They reached “substantial” intercoder agreement (alpha= 0.72) as measured by Krippendorf’s Alpha [47] and met to resolve any disagreements. They collectively finalized the codebook and then divided and independently coded the transcripts. Throughout the paper we provide a relative sense of the frequency of the themes we coded (e.g., that a theme was mentioned by half of participants) but emphasize the importance of not reducing “sensitivity to complex concepts and nuances in data” by viewing these data through a quantitative lens [53].
@@
Ethics. This work was approved by our institution’s ethics review board. We took great care to protect the privacy of participants, which we detail in Appendix B.2. Since 42% of the participants had experienced NDII, we warned all participants that discussing their experiences may be triggering and phrased questions to avoid re-traumatization. To ensure both participant and researcher protection, our team includes a clinical psychologist who reviewed the interview protocols, analyzed data, and supported the research team. We additionally asked our partner organization, the European Sex Workers Rights Alliance (ESWA), and a victim-survivor of NDII for feedback on the protocol. We followed trauma-informed best practices set forth by sexual violence researchers [11] that minimize risk to participants and center victim-survivors’ choice, well-being, empowerment, and autonomy. Team members were trained in response techniques by a licensed clinical psychologist who specializes in sexual assault and trauma (e.g., offering to pause/stop the interview) and were prepared with resources (e.g., hotline information) to offer participants if needed. We note that sharing these experiences may also be a form of empowerment for victim-survivors [32].
@@
Impacts and Research Justice. The interviews were transcribed by a sex worker (trained in transcription and compensated) as part of a research justice commitment to materially support communities we study. We published a white paper of this work through ESWA. We will share published research with participants who requested it and plan to develop a resource guide based on this work. In addition, we have been using these findings to advocate for platform changes and policies that may mitigate IBSA. We translated design recommendations from our participants into wireframes that we have presented to relevant platform product teams. We have also presented findings from this research to policymakers in the European Union, the United States, and Australia. Our partner organization, ESWA, has been using the white paper to conduct workshops on IBSA.
@@
Positionality. This work was conducted by a team of researchers who collectively are scholars of technology, sex work, and sexual violence. This research focuses on adults living in Europe but only half of the research team is based in Europe. Given the diversity of our participants, our team does not completely reflect all of the genders, ethnicities, and sexual orientations of our participants. Therefore, there may have been some limitations in our interpretations of the data.
@@
Limitations. Since we conducted interviews in English and it is not the primary language in many European countries, this may have limited or skewed participation to those in countries where English is more prevalent and/or to those who are more educated, who tend to have studied English. Although we reassured participants that we were interested in understanding their perspectives to advocate for safer technologies for intimate content and took care to phrase questions non-judgmentally, the existing stigma of sharing intimate content may have prevented participants from fully disclosing their behaviors. Since there was a set of features we specifically asked participants about (Appendix B.3), there may be more data on these features; however, each feature in this list was also independently mentioned by participants.
Participant Labeling. We intentionally did not distinguish between quotes from different participant groups and labeled all participants “P#.” Given the stigmatization of sex work, we did not want quotes from sex workers to be dismissed or to be attributed as only relevant toward sex work (and we also asked sex workers about how they share intimate content in their personal lives). In general, there were many commonalities in the narratives shared across all participant groups and we offer Section 8.4 for comparisons between them.
@@
Platform Blinding. With the exception of a few quotes, we replaced specific platforms with the type of platform mentioned in italics. A quote that says, “I use WhatsApp” would be replaced as, “I use [messaging app].” This is to draw focus away from specific platforms and emphasize that intimate content is likely being shared across any and all platforms that enable the storing and/or sharing of images or videos.
3.1 Demographics
When recruiting participants, we considered diversity across different dimensions. We present summaries of participant responses to various pre-interview survey questions below.
Geography. Participants resided in 14 countries: Belgium, France, Germany, Greece, Ireland, Latvia, North Macedonia, Norway, Poland, Portugal, Serbia, Spain, The Netherlands, and the United Kingdom.
Race. Participants were given the option to self-describe their race through text-entry. The majority of participants (65%) identified as Caucasian/White (e.g., “Caucasian,” “White, Russian”) including 2 mixed race individuals. Among those who did not (35%), 15 opted to self-describe their race: 3 participants identified as Black (e.g., “Black British,” “Black African”), 3 identified as Latina/Latino/Latinx, 2 identified as mixed race, and 2 identified as “Brown” or “Southeast Asia, Brown.” Other responses included “French-Moroccan,” “…born in South America,” “Eurasian.”
Education. A majority (73%) of participants had completed a Bachelor’s degree or higher education. 17% had completed some college, 6% had completed high school (or equivalent), and 4% did not finish high school.
Disability. 56% of participants responded that they had a disability (physical, mental health, or neurodivergence) while 35% did not (10% did not answer).
Age. Participants spanned the birth years of 2000 to 1976 with a median birth year of 1994 (roughly 28 years old at time of interview).
Gender & Sexuality. Participants were asked to optionally self-describe their gender and sexual orientation. To present summary statistics, we categorized the free-form text responses we received. These categories therefore represent a close approximation for some responses and do not capture the full richness of responses.
How would you describe your sexual orientation? %
Asexual/Bisexual/Gay/Lesbian/Pansexual/Queer 71%
Heterosexual 21%
Prefer Not to Answer 10%
How would you describe your gender? %
Cis or Trans Woman 27%
Cis or Trans Man 38%
Non-Binary, Gender Fluid, Agender 19%
Prefer Not to Answer 13%
Other 2%
@@ 4 Summary of Findings
We conducted an interview study with 52 people in Europe who share intimate content (24 sex workers, 28 recreational sharers), 42% of whom were victim-survivors of IBSA. Our participants used 40+ technologies to share intimate content. They relied on a combination of heuristics to choose technologies for sharing but nearly half chose to use their default method of communication. Participants’ threat models around NDII were focused on a recipient resharing their content but they were also concerned about data breaches, hacking, and unintentional content discovery. Echoing prior research, all emphasized that stigma exacerbated these threats. To protect themselves, participants implemented a maze of safety strategies that included technical features (e.g., expiring messages) and interpersonal strategies (e.g., setting rules). In sections 5 through 8, we detail sharing contexts, threat models, how participants chose technologies for sharing content, and defensive strategies. In Section 9, we include narratives on how technology could be improved to support digital intimacy/reduce NDII alongside recommendations for platforms and researchers.